# Exciton self-trapping in twisted hexagonal boron nitride homostructures

### v1

@@
The luminescence signal at 300 nm from self-trapped excitons at the interface is likely to be related to a similar signal previously observed in some defected hBN samples [79] as well as in other sp2-hydridized BN samples that exhibit interface disorder, such as turbostratic BN, pyrolytic BN, or BN multiwall nanotubes [80, 81]. Also broad, centered around 4 eV and with long decay dynamics, this emission appears only above 6 eV excitation, i.e. above the excitonic bandgap, suggesting its excitonic origin and a 2 eV energy shift. This emission could also be related to exciton self-trapping, which would then appear to be a fairly common phenomenon in these materials, in the native stacking defect of hBN and more generally in sp2 BN.
@@
Recent studies report unusual exciton-lattice interactions at other h2D interfaces [82, 83, 84, 85], which in some cases could lead to the formation of self-trapped excitons [85]. Our results added to these studies show that the properties of these self-trapped states are easily modulated by the structural properties of h2Ds, such as the twist angle, the choice of stacking, or the thickness of the crystals. h2Ds appear to be an ideal platform for studying and manipulating self-trapped states, opening up new fields of application for these materials. In particular, the high efficiency and the large broadening of the 300 nm luminescence of self-trapped excitons at twisted hBN-hBN interfaces could be exploited for broadband UV light sources [86, 87, 88, 89]. Finally, the formation of interlayer excitons at the interfaces between two TMDs, (with hole and electron spatially separated in the two layers), has been shown to increase the exciton lifetime by more than 3 orders of magnitude [7, 8, 9, 10]. The long lifetime and dark character of the X2D excitons at the hBN-hBN interface might point towards interlayer excitons.
.8 VIII - Conclusion
@@
Sixteen twisted hBN-hBN homostructures have been studied using CL and TRCL as function of temperature with the aim of understanding the broad luminescence observed at a wavelength around 300 nm in hBN-hBN homostructures. Our study shows that this luminescence is indirectly excited by the transport of free excitons from the volume of the flakes to the interface via their diffusion. We have shown that the internal quantum yield of the luminescence at 300 nm is close to 100% at angles close to 30° on HPHT samples. This luminescence is attributed to the radiative recombination of self-trapped excitons at the interface. The exciton dipole induces a local deformation of the interface lattice, in which the exciton is ”self-trapped”.
@@
TRCL luminescence decays could be analyzed by considering a trapping-detrapping model where 3 exciton populations are coupled: a population of free excitons in the volume (X3D), a population of excitons trapped at the 2D interface (X2D, not self-trapped), and a population of excitons self-trapped at the interface (XST). In this scenario, which is consistent with our experimental data set, excitons accumulate at the interface in a non-radiative (dark) and long living X2D form. Analysis of the data as a function of temperature revealed a self-trapping energy barrier of the order of 10 meV and an interface trapping potential for excitons of about 100 meV on two samples with twist angles of 11 and 15°. These values agree with the literature and theoretical calculations available to date.
@@
For highly twisted hBN-hBN structures, the ionic character and the deformability of the lattice are locally and densely perturbed across the entire interface, which is probably at the root of exciton self-trapping. Beyond these qualitative elements, it remains to be understood in detail what is the key element that causes self-trapping only at high angles, and further theoretical studies would be welcome. Moreover, the physical phenomena occurring in h2Ds at high angles are very different from those at small angles with Moiré superstructures. To explore them, one avenue would be to apply the quasi-crystal physics to h2Ds with high twist angles.
@@ Acknowledgements.
Appendix A : Measurement of the internal quantum efficiency of the broad 300 nm luminescence.
The procedure to measure the internal quantum efficiency ($IQE$) of a luminescence band during cw-CL experiment is described in ref. [51]. In this reference, it is applied to measure the $IQE$ of the free exciton luminescence of hBN single crystals that appears at 215 nm. To apply the same procedure to the 300 nm band, one should consider the significative change of the refractive index of the top surface between 215 and 300 nm.
In our case, the CL intensity is integrated between 240 and 400 nm as shown in Fig. 9. The reflection index $n$ of the surface is decreases from $n$=3 to $n$=2.37 over the 240-400 nm range [Segura2018], corresponding to a light extraction that varies from 2.1% to 3.9%. To simply measure the magnitude of the $IQE$ of this band, we approximate a constant index $n$=2.56 measured at the band maximum (3.2% extraction). This approximation is reasonable given the uncertainty of the $IQE$ measurement, which is of the order of 50% [51].
Figure 9: Corrected CL spectrum of a hBN-hBN homostructure at 300 K. The integrated intensity of the free exciton band of hBN appears in red, and the integrated intensity of the broad optical emission at 300 nm, characteristic of twisted hBN-hBN homostructures, appears in blue.
@@
Fig. 10 shows spectra measured on a slightly twisted PDC sample in green and a highly twisted PDC sample in red, compared to a spectrum measured on a single PDC crystal in black. The single crystal shows emission from point defects (labeled $\alpha$, $\beta$, $\gamma$) that occur locally in some crystals. As shown in Fig. 4(c) of the main text, the CL intensity of the hBN-hBN interface decreases with decreasing twist angle. For the PDC-2° sample, the intensity of the interface luminescence is lower than the intensity of the native color center emission. To avoid confusion between native defects and interface emission, only highly twisted single crystal structures without significant native defect emission in the 200-400 nm range have been investigated.
@@
Appendix D : Luminescence decay as function of the energy within the broad 300 nm luminescence.
Given the linewidth of the emission at 300 nm, which extends between 250 and 400 nm (i.e., nearly 2 eV), the question arises whether it is the result of a population of emitters of different energies or whether the signal is intrinsically broad. To check this, the decay of the luminescence is studied at different energies of the broad emission.
Fig. 11(a) shows the spectrum of the APHT-28° sample, dominated by the luminescence of interest with little Fabry-Pérot interference contrast. Fig. 11(b) shows the TRCL decays of the filtered luminescence at different wavelengths ($\pm$ 7.5 nm). The decay dynamics are perfectly identical for all investigated wavelengths. This result suggests a unique recombination mechanism for this particularly broad band. The large spectral width of the 300 nm band thus appears to be an intrinsic feature of the luminescence process of twisted hBN-hBN homostructures.
Figure 11: CL spectrum (3 kV, 280 pA) measured on the APHT-28° homostructure at 300 K. (b) Decay of the luminescence intensity, filtered at different wavelengths, recorded in TRCL after the excitation was stopped at t = 0. The wavelengths corresponding to each color are indicated by arrows in (a).
@@
Fig. 12 shows spectra obtained on a 23° twisted hBN-hBN structure and on a hBN single crystal at room temperature and at cryogenic temperature. On the single crystal, the luminescence is dominated by the free exciton luminescence occurring at 215 nm, while on the hBN-hBN structure the emission is dominated by the broad luminescence occurring between 250 and 400 nm. The band is modulated by Fabry-Perot interference. We observe that the band remains extremely broad even at cryogenic temperatures. A color center emission also appears on the hBN-hBN structure and on the single hBN crystals, with a zero-phonon band at 305 nm ($\alpha$) and its phonon replica ($\beta$ and $\gamma$). This emission shows a large thermal broadening typical of color centers, indicating that its nature is different from that of the hBN-hBN interface emission, whose broadening remains around 2 eV even when the sample holder temperature reaches 5K.
@@
The TRCL decay of the luminescence intensity is related to the evolution of the exciton populations and thus to their interactions. The scenario of exciton population interactions discussed in the main text is the following: the free exciton population NX3D, generated in the crystal volume by the primary excitation, are trapped (rate $C$) at the interface. The excitons accumulate at the 2D interface under a dark and long living form (NX2D population). They are finally converted into NXST by their self-trapping (rate $ST$) and into NX3D by their detrapping (rate $D$). This scenario, illustrated in Fig. 7 of the main text, leads to the following differential equations:
$\left\{\begin{array}[]{rcl}\frac{dN_{X3D}}{dt}&=&-\frac{N_{X3D}}{\tau_{0}}-N_{X3D}\times C+N_{X2D}\times D+G(t)\\ \\ \frac{dN_{X2D}}{dt}&=&-N_{X2D}(D+ST)+N_{X3D}\times C\\ \\ \frac{dN_{XST}}{dt}&=&-\frac{N_{XST}}{\tau_{XST}}+N_{X2D}\times ST\\ \end{array}\right.$ (1)
where $\tau_{0}$ is the lifetime of the free exciton in the volume (without interface) and $\tau_{XST}$ is the lifetime of the XST excitons. $G$ denotes the generation rate of free excitons, which is constant during continuous excitation and then zero when the excitation stops. A priori, $C$ varies slightly with temperature due to the weak evolution of the exciton diffusion length with temperature [61], while the rates of $D$ and $ST$ are thermally activated.
The first two equations of the system are not coupled to NXST, they could be considered independently. We find a superposition of two exponential decays for both NX3D and NX2D exciton populations: one short with a characteristic time $\tau_{s}$ and the other long with a characteristic time $\tau_{l}$ $>$ $\tau_{s}$. These times can be written with A = 1/$\tau_{0}$ + $C$, and B = $D$ + $ST$:
@@
$\frac{N_{X2D}(0)}{N_{X3D}(0)}=\frac{C}{B}$ (4)
And the NX3D population is written as:
$\frac{N_{X3D}(t)}{N_{X3D}(0)}=A_{l}e^{-\frac{t}{\tau_{l}}}+(1-A_{l})e^{-\frac{t}{\tau_{s}}}$ (5)
@@
The parameter B=$D$+$ST$ can therefore be extracted from the TRCL decays of the free exciton luminescence from the previous equation, or from the simplified form when Al <<much-less-than$<$ < 1 :
$D+ST\approx\frac{1}{\tau_{l}}$ (8)
Calculations show that all three populations follow the same decay time, as can be seen in Fig. 13(a), which is consistent with the experimentally observed decays of the NXST and NX3D populations in Fig. 13(a) (see also Fig. 3(b) in the main text). The long component of the TRCL decay being governed by the detrapping and self-trapping rates of the NX2D population that occurs in the long term.
Figure 13: (a) Theoretical decays of the NX2D and NX3D exciton populations for $\tau_{s}$=10 ns, $\tau_{l}$=100 ns and Al=0.1, and of the NXST self-trapped exciton population for $\tau_{XST}$ = 10 ns. (b) Experimental decays for the NXST (at 300 nm) and NX3D (at 215 nm) exciton populations on the HPHT-15° structure measured in TRCL at 5 kV, 27 pA, 300 K.
@@
  • Li et al. [2018] Y. Li, V. Garnier, C. Journet, J. Barjon, A. Loiseau, I. Stenger, A. Plaud, B. Toury, and P. Steyer, Advanced synthesis of highly crystallized hexagonal boron nitride by coupling polymer-derived ceramics and spark plasma sintering processes—influence of the crystallization promoter and sintering temperature, Nanotechnology 30, 035604 (2018).
@@
  • Roux et al. [2023] S. Roux, C. Arnold, E. Carré, E. Janzen, J. H. Edgard, C. Maestre, B. Toury, C. Journet, V. Garnier, P. Steyer, et al., Surface recombination and out of plane diffusivity of free excitons in hexagonal boron nitride, arXiv preprint arXiv:2308.05539  (2023).
@@
  • Paleari et al. [2019] F. Paleari, H. P. Miranda, A. Molina-Sánchez, and L. Wirtz, Exciton-phonon coupling in the ultraviolet absorption and emission spectra of bulk hexagonal boron nitride, Physical review letters 122, 187401 (2019).
  • Paleari [2019] F. Paleari, First-principles approaches to the description of indirect absorption and luminescence spectroscopy: exciton-phonon coupling in hexagonal boron nitridePh.D. thesis (2019).

### v2

@@
This work on self-trapped excitons at twisted hBN-hBN interfaces might cast a new light on similar luminescence signal reported in defective hBN samples [84] as well as in other sp2-hydridized BN samples that exhibit interface disorder, such as turbostratic BN, pyrolytic BN, or BN multiwall nanotubes [85, 86]. Also typically 2 eV broad, centered around 4 eV and with long decay dynamics, this emission is excited only above 6 eV excitation, i.e. above the excitonic bandgap, suggesting an excitonic origin. This similar emission observed in disordered BN materials could also be related to exciton self-trapping, which would then appear to be a fairly common phenomenon in the BN material familly.
@@
Recent studies report unusual exciton-lattice interactions at other h2D interfaces [87, 88, 89, 90], which in some cases could lead to the formation of self-trapped excitons [90]. Our results added to these studies show that the properties of these self-trapped states are easily modulated by the structural properties of h2Ds, such as the twist angle, the choice of stacking, or the thickness of the crystals. h2Ds appear to be an ideal platform for studying and manipulating self-trapped states, opening up new fields of application for these materials. In particular, the high efficiency and the large broadening of the 4-eV luminescence of self-trapped excitons at twisted hBN-hBN interfaces could be exploited for broadband UV light sources [91, 92, 93, 94].
Finally, the formation of interlayer excitons at the interfaces between two TMDs, (with hole and electron spatially separated in the two layers), has been shown to increase the exciton lifetime by more than 3 orders of magnitude [7, 8, 9, 10]. The long lifetime and dark character of the X2D excitons at the hBN-hBN interface might have common features with interlayer excitons.
.9 IX - Conclusion
@@
Sixteen twisted hBN-hBN homostructures have been studied using CL and TRCL as function of temperature with the aim of understanding the broad 4-eV luminescence in hBN-hBN homostructures. Our study shows that this luminescence is indirectly excited by the transport of free excitons from the upper and lower hBN flakes to the interface between them via their diffusion and capture. Accumulated at the interface, excitons probably undergo exciton-exciton annihilation which manifests by the decrease of the light emission efficiency at moderate excitation powers. In the low excitation regime, it was found that the efficiency of the 4-eV emission reaches almost 100% at angles close to 30°. This intense light emission is attributed to the radiative recombination of self-trapped excitons at the interface. A reduction of the excitons size after their capture at the interface is believed to favor a local deformation of the crystal around them in which they self-trap.
@@
TRCL luminescence decays could be analyzed by considering a trapping-detrapping phenomenological model where 3 exciton populations are coupled: free excitons in the volume (X3D), excitons trapped at the 2D interface (X2D, not self-trapped), and excitons self-trapped at the interface (XST). In this scenario consistent with our set of experimental data, excitons accumulate at the interface in the non-radiative (dark) and long living X2D form. Our analysis of temperature-dependant experiments for twist angles of $\sim$15° reveals a self-trapping energy barrier of a few 10 meV and an interface trapping potential measured around 100 meV. The quantitative analysis of TRCL data further provides an estimation of the detrapping rate near 15° and of the self-trapping rate near 30°. They suggest a deeper interface potential near 30° twist angles, limiting exciton detrapping from the interface and thus favoring self-trapping. This is consistent with the remarkably high luminescence efficiency of the XST luminescence at high twist angles. These values agree with the literature and theoretical calculations available to date.
@@
For highly twisted hBN-hBN structures, the crystal lattice appears to be locally and densely perturbed across the entire interface. Beyond these qualitative elements, it remains to be understood in detail what is the key elements that cause self-trapping. The physical phenomena occurring in h2Ds at high angles appear very different from those at small angles with moiré superstructures. To explore them, one avenue would be to apply the quasi-crystal physics to h2Ds with high twist angles.
@@ Acknowledgements.
Appendix A: Measurement of the internal quantum efficiency of the 4-eV luminescence.
The procedure to measure the internal quantum efficiency ($IQE$) of a luminescence band during CL experiment under continuous excitation is described in ref. [51]. In this reference, it is applied to measure the $IQE$ of the free exciton luminescence of hBN single crystals that appears at 5.75 eV. To apply the same procedure to the 4-eV band, one should consider the significative change of the refractive index of the top surface between 5.75 and 4-eV.
In our case, the CL intensity is integrated between 3.1 and 5.2 eV as shown in Fig. 11. The reflection index $n$ of the surface is increased from $n$=2.37 to $n$=3 over the 3.1-5.2 eV range [95], corresponding to a light extraction that varies from 3.9% to 2.1%. To simply measure the magnitude of the $IQE$ of this band, we approximate a constant index $n$=2.56 measured at the band maximum (3.2% extraction). This approximation is reasonable given the uncertainty of the $IQE$ measurement, which is of the order of 50% [51].
Figure 11: Corrected CL spectrum of a hBN-hBN homostructure at 300 K. The integrated intensity of the free exciton band of hBN appears in red, and the integrated intensity of the 4-eV luminescence, characteristic of twisted hBN-hBN homostructures, appears in blue.
@@
Figure 12 shows spectra measured on a slightly twisted PDC sample in green and a highly twisted PDC sample in red, compared to a spectrum measured on a single PDC crystal in black. The single crystal shows emission from point defects (labeled $\alpha$, $\beta$, $\gamma$) that occur locally in some crystals. As shown in Fig. 5(c) of the main text, the CL intensity of the hBN-hBN interface decreases with decreasing twist angle. For the PDC-2° sample, the intensity of the interface luminescence is lower than the intensity of the native color center emission. To avoid confusion between native defects and interface emission, only highly twisted single crystal structures without significant native defect emission in the 3-6 eV range have been investigated.
@@
Appendix D: Luminescence decay as function of the energy within the broad 4-eVm luminescence.
Given the linewidth of the emission at 4-eV luminescence, which extends between 3 and 5 eV, the question arises whether it is the result of a population of emitters of different energies or whether the signal is intrinsically broad. To check this, the decay of the luminescence is studied at different energies of the broad emission.
Figure 13(a) shows the spectrum of the APHT-29° sample, dominated by the luminescence of interest with little Fabry-Pérot interference contrast. Figure 13(b) shows the TRCL decays of the luminescence filtered at different energies ($\pm$0.15 eV). The decay dynamics are perfectly identical for all investigated energies. This result suggests a unique recombination mechanism for this particularly broad band. The large spectral width of the 4-eV band thus appears to be an intrinsic feature of the luminescence process of twisted hBN-hBN homostructures.
Figure 13: CL spectrum (3 kV, 280 pA) measured on the APHT-29° homostructure at 300 K. (b) Decay of the luminescence intensity, filtered at different energies, recorded in TRCL after the excitation was stopped at t = 0. The energies corresponding to each color are indicated by arrows in (a).
@@
Figure 14 shows spectra obtained on a 23° twisted hBN-hBN structure and on a hBN single crystal at room temperature and at cryogenic temperature. On the single crystal, the luminescence is dominated by the free exciton luminescence occurring at 5.75 eV, while on the hBN-hBN structure the emission is dominated by the broad luminescence occurring between 3 and 5 eV. The band is modulated by Fabry-Perot interference. We observe that the band remains extremely broad even at cryogenic temperatures. A color center emission also appears on the hBN-hBN structure and on the single hBN crystals, with a zero-phonon band at 4.06 eV ($\alpha$) and its phonon replica ($\beta$ and $\gamma$). This emission shows a large thermal broadening typical of color centers, indicating that its nature is different from that of the hBN-hBN interface emission, whose broadening remains around 2 eV even when the sample holder temperature reaches 5K.
@@
The TRCL decay of the luminescence intensity is related to the evolution of the exciton populations and thus to their interactions. The scenario of exciton population interactions discussed in the main text is the following: the free exciton population $N_{3D}$, generated in the crystal volume by the primary excitation, are trapped (rate $C$) at the interface. The excitons accumulate at the 2D interface under a dark and long living form ($N_{2D}$ population). They are finally converted into NST by their self-trapping (rate $ST$) and into N3D by their detrapping (rate $D$). This scenario, illustrated in Fig. 8 of the main text, leads to the following differential equations:
$\left\{\begin{array}[]{rcl}\frac{dN_{3D}}{dt}&=&-\frac{N_{3D}}{\tau_{0}}-CN_{3D}+DN_{2D}+G(t)\\ \\ \frac{dN_{2D}}{dt}&=&-(D+ST)N_{2D}+CN_{3D}\\ \\ \frac{dN_{ST}}{dt}&=&-\frac{N_{ST}}{\tau_{ST}}+STN_{2D}\\ \end{array}\right.$ (1)
where $\tau_{0}$ is the lifetime of the free exciton in the volume (without interface) and $\tau_{ST}$ is the lifetime of the XST excitons. $G$ denotes the generation rate of free excitons, which is constant during continuous excitation and then zero when the excitation stops. A priori, $C$ varies slightly with temperature due to the weak evolution of the exciton diffusion length with temperature [63], while the rates of $D$ and $ST$ are thermally activated.
The first two equations of the system are not coupled to NST, they could be considered independently. We find a superposition of two exponential decays for both N3D and N2D exciton populations: one short with a characteristic time $\tau_{s}$ and the other long with a characteristic time $\tau_{l}$ $>$ $\tau_{s}$. These times can be written with A = 1/$\tau_{0}$ + $C$, and B = $D$ + $ST$:
@@
$\frac{N_{2D}(0)}{N_{3D}(0)}=\frac{C}{B}$ (4)
And the N3D population is written as:
$\frac{N_{3D}(t)}{N_{3D}(0)}=A_{l}e^{-\frac{t}{\tau_{l}}}+(1-A_{l})e^{-\frac{t}{\tau_{s}}}$ (5)
@@
The parameter B=$D$+$ST$ can therefore be extracted from the TRCL decays of the free exciton luminescence from the previous equation, or from the simplified form when Al $\ll$ 1:
$\frac{1}{\tau_{l}\simeq D+ST}$ (8)
Calculations show that all three populations follow the same decay time, as can be seen in Fig. 15(a), which is consistent with the experimentally observed decays of the NST and N3D populations in Fig. 15(a) (see also Fig. 4(b) in the main text). The long component of the TRCL decay being governed by the detrapping and self-trapping rates of the N2D population that occurs in the long term.
Figure 15: (a) Simulated decays of the N2D and N3D exciton populations for $\tau_{s}$=10 ns, $\tau_{l}$=100 ns and Al=0.1, and of the NST self-trapped exciton population for $\tau_{ST}$ = 10 ns. (b) Experimental decays for the NST and N3D exciton populations on the HPHT-15° structure measured in TRCL at 5 kV, 27 pA, 300 K.
@@
  • Hoffman et al. [2014] T. B. Hoffman, B. Clubine, Y. Zhang, K. Snow, and J. H. Edgar, Optimization of ni–cr flux growth for hexagonal boron nitride single crystals, Journal of crystal growth 393, 114 (2014).
@@
  • Maestre et al. [2022] C. Maestre, Y. Li, V. Garnier, P. Steyer, S. Roux, A. Plaud, A. Loiseau, J. Barjon, L. Ren, C. Robert, et al., From the synthesis of hbn crystals to their use as nanosheets in van der waals heterostructures, 2D Materials 9, 035008 (2022).
  • Ouaj et al. [2024] T. Ouaj, C. Arnold, J. Azpeitia, S. Baltic, J. Barjon, J. Cascales, H. Cun, D. Esteban, M. Garcia-Hernandez, V. Garnier, et al., Benchmarking the integration of hexagonal boron nitride crystals and thin films into graphene-based van der waals heterostructures, arXiv preprint arXiv:2409.03652  (2024).
@@
  • Roux et al. [2024] S. Roux, C. Arnold, E. Carré, E. Janzen, J. H. Edgar, C. Maestre, B. Toury, C. Journet, V. Garnier, P. Steyer, et al., Surface recombination and out-of-plane diffusivity of free excitons in hexagonal boron nitride, Physical Review B 109, 155305 (2024).
  • Plaud et al. [2019] A. Plaud, L. Schué, K. Watanabe, T. Taniguchi, F. Fossard, F. Ducastelle, A. Loiseau, and J. Barjon, Exciton-exciton annihilation in hbn, Applied Physics Letters 114 (2019).
  • Note [1] Experiments are performed at 10 K in ref. [64], and at 300 K in ours. However, since in such samples the diffusivity is temperature independent [63], we do not expect drastic differences in the probability of exciton-exciton collisions and thus in the EEA rate as the temperature is increased.
@@
  • Fournier et al. [2021] C. Fournier, A. Plaud, S. Roux, A. Pierret, M. Rosticher, K. Watanabe, T. Taniguchi, S. Buil, X. Quélin, J. Barjon, et al., Position-controlled quantum emitters with reproducible emission wavelength in hexagonal boron nitride, Nature communications 12, 3779 (2021).
@@
  • Paleari et al. [2018] F. Paleari, T. Galvani, H. Amara, F. Ducastelle, A. Molina-Sánchez, and L. Wirtz, Excitons in few-layer hexagonal boron nitride: Davydov splitting and surface localization, 2D Materials 5, 045017 (2018).
  • Galvani et al. [2016] T. Galvani, F. Paleari, H. P. Miranda, A. Molina-Sánchez, L. Wirtz, S. Latil, H. Amara, and F. Ducastelle, Excitons in boron nitride single layer, Physical Review B 94, 125303 (2016).
@@
  • Segura et al. [2018] A. Segura, L. Artús, R. Cuscó, T. Taniguchi, G. Cassabois, and B. Gil, Natural optical anisotropy of h-bn: Highest giant birefringence in a bulk crystal through the mid-infrared to ultraviolet range, Physical Review Materials 2, 024001 (2018).